ln(log(x)-2)=5 Solve for x.
ln(a+b)+ln(a-b)-2lnc Use rules of logs to combine the expression.

Answer by Alan3354(69443)

ln(log(x)-2)=5 Solve for x.
--------------
log(x)-2 = e^5
log(x) = e^5 + 2
x = 10^(e^5+2)
x =~ 10^150.413159 (= a big number)